# Models, Methods and Waveforms for Estimation and Prediction of Doubly Sparse Time-Varying Channels

### v1

@@ III-C Compressed-sensing estimation of DS-LTV channels using different waveforms
Thanks to Theorems 1 and 2, matrix $\mathbf{M}_{\rm p}$ with large enough $L,Q,N_{\rm p}$ can be made to have a HiRIP constant that satisfies $\delta_{3s_{\rm d},2s_{\rm D}}<\frac{1}{\sqrt{3}}$. The conditions of [18, Theorem 1] are thus satisfied, and the corollary follows from that theorem. ∎
To get some insights into the above results, we look at them in an asymptotic regime defined with the help of an auxiliary variable $K$: $L\sim K$, $Q\sim K$ ($L,Q\sim K$ implies $N\sim K^{2}$ [16]), $s_{\rm d}\sim K^{\kappa_{\rm d}}$ and $s_{\rm D}\sim K^{\kappa_{\rm D}}$ for some $\kappa_{\rm d},\kappa_{\rm D}\in[0,1)$. In this regime, SCM pilot and guard overhead satisfies $(2L-1)N_{\rm p,t}=\Omega\left((\log K)^{2}K^{1+\kappa_{\rm D}}\right)$ due to Theorem 1. The pilot, guard and CP overhead of OFDM equals $\left(L-1+N_{\rm p,f}\right)N_{\rm p,t}$, and is thus smaller than its SCM counterpart but has the same asymptotic growth. As for AFDM, pilot and guard overhead satisfies $N_{\rm p}\left((L-1)P+2Q+1\right)=\Omega\left(\log(\log K)(\log K)^{2}K\right)$. Therefore, under double delay-Doppler sparsity, AFDM overhead is asymptotically dominated by that of OFDM and SCM.
The next section is dedicated to a relaxation of the on-grid approximation to generalize the above results to more realistic propagation and to improve compatibility with channel prediction.
IV Estimation and extrapolation of doubly sparse linear time-varying channels with off-grid Doppler shifts
While the on-grid approximation of Definition 1 is useful for the conception of LTV channel estimation and sensing schemes and for the analysis of their performance as we argued in Section III, it lacks support for the finer Doppler resolution needed for channel prediction or mitigation of channel aging. Indeed, most channel prediction paradigms [20, 21, 22, 23] involve, explicitly of implicitly, the estimation of Doppler frequency shifts to within an error margin smaller than the frequency resolution characteristic of the duration of the channel observation interval. For that sake, we now present a second approximation for LTV channels that allows, in contrast to the first approximation in Definition 1, for fractional-valued Doppler frequency shifts.
@@ IV-A Off-grid doubly sparse linear time-varying channels
We now allow for fractional-valued Doppler frequency shifts. In this case, only $\iota_{p}$ is zero, while $\kappa_{p}$ may take non-zero values sampled from a uniform distribution i.e., $\kappa_{p}\sim\mathcal{U}([-\frac{1}{2},\frac{1}{2}])$. Define $N_{\mathrm{D},l,q}$ as the number of “sub”-paths with the same delays and Doppler shifts integer part and which only differ in their Doppler shifts fractional part:
@@
In this configuration, the number of paths $N_{\rm p}$ in (2) satisfies $N_{\rm p}=\sum_{l=0}^{L-1}\sum_{q=-Q}^{Q}N_{\mathrm{D},l,q}I_{l,q}$. Depending on the scenario, it is in principle possible to model $N_{\mathrm{D},l,q}$ either as a fixed value or as a random variable, for example, a uniform random variable drawn over the range $\llbracket 1\mkern 1.5mu..\mkern 1.5muN_{\rm D}\rrbracket$. For simplicity, we opt for the first option, that is, $N_{\mathrm{D},l,q}=N_{\rm D},\forall l,q$ for some value $N_{\rm D}$ and we impose that the following channel power normalization should be satisfied.
@@
Definition 4 (off-grid Delay-Doppler double sparsity).
The complex gain $h_{l,n}$ of the $l$-th path of the LTV channel varies with time as
@@
for some value $N_{\rm D}$, with $I_{l,q}$ retaining the same description provided in Definition 1 and with $I_{l,q},\left\{\alpha_{l,q,i},\kappa_{i}\right\}_{i=1\cdots N_{\rm D}}$ being mutually independent. The complex gains are assumed to satisfy $\alpha_{l,q,i}\sim\mathcal{CN}\left(0,\sigma_{\alpha}^{2}\right)$ with $\sigma_{\alpha}^{2}$ chosen so that (18) is respected.
@@
Maintaining the delay taps as integer values is not to the detriment of the core message of this work since the same approach we develop to deal with off-grid Doppler shifts can be extended to off-grid delays. Moreover, the effect of off-grid delays can still be made to fall under the current model by increasing the value of the model delay domain sparsity parameter $p_{\rm d}$ sufficiently to account for leakage due to the fractional part of the delay shifts.
@@
Estimating all the unknown parameters of the off-grid channel model in (19) remains a challenging task. Indeed, the $N_{\rm D}$ Doppler frequency shifts $\{q+\kappa_{i}\}_{i=1\cdots N_{\rm D}}$ associated with a grid point $(l,q)$ are only different by their fractional part and hence render the problem of estimating the corresponding complex coefficients $\{\alpha_{l,q,i}\}_{i=1\cdots N_{\rm D}}$ ill-conditioned. Moreover, the number of unknowns in the model i.e., $N_{\rm p}$, could be prohibitively large. For both these reasons, we propose a new model based on multiple “elementary” basis expansion models (BEM), each set with a bandwidth equal to the frequency resolution and shifted in frequency to be centered at one of the active grid points. Thanks to its optimality in terms of time-frequency localization [24], we opt for BEM based on discrete prolate spheroidal sequences (DPSS).
@@ IV-B Background: DPSS Basis Expansion Model
We begin with an overview of DPSS BEM modeling of a baseband signal $h_{n}$ that occupies a bandwidth $(-W,W)$ in digital frequencies with $W\in\left(0,\frac{1}{2}\right)$. DPSS basis vectors $u_{b,n}^{(N,W)}$ ($b=1,\ldots,N$) are the eigenvectors of the prolate matrix [25]:
$\sum_{k=0}^{N-1}C_{k,n}^{(N,W)}u_{b,k}^{(N,W)}=\lambda_{b}^{(N,W)}u_{b,n}^{(N,W)},b=1\cdots N,n=0\cdots N-1$ (20)
where $C_{k,n}^{(N,W)}$ is the $(k-n)$-th entry of the prolate matrix:
@@
The eigenvectors are normalized so that $\sum_{n=1}^{N}\left(u_{b,n}^{(N,W)}\right)^{2}=1$. The eigenvalues are ordered according to their values starting with the largest one: $1\geq\lambda_{0}^{(N,W)}\geq\ldots\geq\lambda_{N-1}^{(N,W)}\geq 0$. The eigenvalues $\lambda_{b}^{(N,W)}$ (representing energy concentration) are clustered near $1$ for $b\leq 2WN$, and rapidly drop to zero for $b>2WN$ [26]. In this study we employ multiple DPSS BEMs, each of which is used to represent the channel signal component related to the fractional part of the $N_{\rm D}$ Doppler shifts around one of the grid points $(l,q)$ i.e., $\sum_{i=1}^{N_{\rm D}}\alpha_{l,q,i}e^{\imath 2\pi\frac{n\kappa_{i}}{N}}$, and not the channel signal associated with the whole Doppler spread i.e., $h_{l,n}$. Each of these BEMs is defined using (20) and (21) with $W=\frac{1}{2N}$. In the interest of simplicity, all parameters with the superscript $(N,W)$ in their notation, such as $u_{b,n}^{(N,W)}$, will be replaced by their simplified forms e.g., $u_{b,n}$.
IV-C Off-grid approximation using multiple shifted elementary BEMs
The following term in (19) represents a signal occupying a frequency band centered at digital frequency $\frac{q}{N}$ and of a bandwidth equal to $\frac{1}{N}$
$h_{l,q,n}\triangleq\sum_{i=1}^{N_{\rm D}}\alpha_{l,q,i}e^{\imath 2\pi\frac{n(q+\kappa_{i})}{N}}$ (22)
Its baseband version defined as $e^{-\imath 2\pi\frac{nq}{N}}h_{l,q,n}$ can thus be modeled (see Figure 6) using a DPSS BEM of an order equal to $Q_{\rm BEM}$ (chosen large enough depending on the required modeling precision) and a bandwidth $(-W,W)$ with $W=\frac{1}{2N}$ by defining
$h_{l,q,n}^{\rm BEM}\triangleq e^{\imath 2\pi\frac{nq}{N}}\sum_{b=1}^{Q_{\rm BEM}}\beta_{l,q,b}u_{b,n}\>.$ (23)
Refer to caption Refer to caption
(a) (b)
Figure 6: (a) Capturing the leakage due to $S_{\rm D}$ clusters of off-grid Doppler shifts $\left\{q_{s}+\kappa_{i}\right\}_{\begin{subarray}{c}s=1\cdots S_{\rm D},\\ i=1\cdots N_{\rm D}\end{subarray}}$ by $S_{\rm D}$ frequency shifted copies of a DPSS BEM $\left\{u_{b,n}\right\}_{b=1\cdots Q_{\rm BEM}}$ ($\tilde{u}_{b,k}$ in the figure being the DFT of $u_{b,n}$). (b) Absolute-value comparison between a DS-LTV channel $h_{l,n}$ having $S_{\rm D}=3$ Doppler shift clusters and its representation using 3 elementary shifted BEMs with $Q_{\rm BEM}=3$ and $Q_{\rm BEM}=4$ ($N=2048$, $W=\frac{1}{2N}$, $Q=7$)
In vector form
@@ IV-C Off-grid approximation using multiple shifted elementary BEMs
with $\mathbf{U}_{Q_{\rm BEM}}\triangleq\left[\mathbf{u}_{1}\ldots\mathbf{u}_{Q_{\rm BEM}}\right]$ and $\mathbf{P}^{\rm BEM}=\mathbf{U}_{Q_{\rm BEM}}\mathbf{U}_{Q_{\rm BEM}}^{\rm H}$ being the orthogonal projection matrix associated with the DPSS BEM representation and $\mathbf{E}_{f}\triangleq\mathrm{diag}\left(e^{\imath 2\pi f0}\ldots e^{\imath 2\pi f(N-1)}\right)$.
Inserting the elementary BEMs defined in (23) for each Doppler grid point into (19) gives
$h_{l,n}^{\rm BEM}$ $=\sum_{q=-Q}^{Q}I_{l,q}e^{\imath 2\pi\frac{nq}{N}}\sum_{b=1}^{Q_{\rm BEM}}\beta_{l,q,b}u_{b,n},$
@@
The precision of the above representation in relation to the order of the elementary BEMs in the model (IV-C) is provided by the following theorem.
@@
For any $\epsilon>0$, if we set $Q_{\rm BEM}\geq C\log\frac{1}{\epsilon}$ for some constant $C$ then the representation error of the model in (IV-C) satisfies $\sum_{l=0}^{L-1}\mathbb{E}\left[\left|h_{l,n}-h_{l,n}^{\rm BEM}\right|^{2}\right]<\epsilon$ for a sufficiently large $N$.
@@
Theorem 3 states that the number of DPSS basis functions needed to represent the channel component associated with a single delay-Doppler grid point grows only logarithmically with the inverse of the target precision. The following figure (as well as the numerical results of Section V) shows that as few as 4 DPSS basis functions ($Q_{\rm BEM}=4$) are sufficient to get good enough precision. Note that this value is unrelated to the Doppler spread value i.e., $12Q+$ 1: a larger Doppler spread does not imply the need for a larger value of $Q_{\rm BEM}$.
IV-D DS-LTV off-grid channel estimation
Let $\boldsymbol{\beta}$ be the vectorized form of the BEM coefficients $\{\beta_{l,q,b}\}_{\begin{subarray}{c}b=1\cdots Q_{\rm BEM}\\ l,q,I_{l,q}\neq 0\end{subarray}}$ associated with the active delay-Doppler grid points. Note that the knowledge of the delay-Doppler sparsity support i.e., of $\left\{I_{l,q}\right\}_{l=0\cdots L-1,q=-Q\cdots Q}$, at the receiver side is now assumed. Define $\mathbf{B}$ as the following block-diagonal matrix
$\mathbf{B}=\mathrm{blkdiag}\left(\widetilde{\mathbf{B}},\ldots,\widetilde{\mathbf{B}}\right)$ (27)
with $\widetilde{\mathbf{B}}\triangleq\left[\widetilde{\mathbf{B}}_{-Q}\ldots\widetilde{\mathbf{B}}_{Q}\right]$ being a $N\times(2Q+1)Q_{\rm BEM}$ matrix satisfying $\left[\widetilde{\mathbf{B}}_{q}\right]_{n,b}=\frac{1}{\sqrt{N}}\tilde{U}_{b,(n-q)}$ and $\tilde{U}_{b,n}=\frac{1}{\sqrt{N}}\sum_{k=0}^{N-1}u_{b,k}e^{-\imath 2\pi\frac{nk}{N}}$ being the DFT of the DPSS basis vector $u_{b,n}$. Let $\mathbf{A}_{\beta}$ be the matrix that places the blocks of $\boldsymbol{\beta}$, each of size $Q_{\rm BEM}$, within the positions corresponding to $I_{l,q}\neq 0$ in a larger vector of length $LQ_{\rm BEM}(2Q+1)$ resulting in a $LQ_{\rm BEM}(2Q+1)$-long vector $\mathbf{A}_{\beta}\boldsymbol{\beta}$ that is block sparse. Now, define

### v2

@@ III-C Theoretical analysis of sparse recovery of DS-LTV channels with different waveforms
Thanks to Theorems 1 and 2, matrix $\mathbf{M}_{\rm p}$ with large enough $L,Q,N_{\rm p}$ can be made to have a HiRIP constant that satisfies $\delta_{3s_{\rm d},2s_{\rm D}}<\frac{1}{\sqrt{3}}$. The conditions of [24, Theorem 1] are thus satisfied, and the corollary follows from that theorem. ∎
Remark 1.
To gain some insights into the above results, we examine them in an asymptotic regime defined with the help of an auxiliary variable $K$. The regime is characterized by $L\sim K$, $Q\sim K$, $s_{\rm d}\sim K^{\kappa_{\rm d}}$ and $s_{\rm D}\sim K^{\kappa_{\rm D}}$ for some $\kappa_{\rm d},\kappa_{\rm D}\in[0,1)$ where the notation $\sim$ stands for asymptotic equivalence. Note that, in this regime, the frame size satisfies $N\sim K^{2}$ due to the fact that $L,Q\sim K$ [21]. As for the fraction of $N$ dedicated to pilot and guard samples (including both prefix and data guard samples), its asymptotic behavior depends on the waveform. Due to Theorem 1, SCM overhead equals
$L-1+(2L-1)N_{\rm p,t}=\Omega\left((\log K)^{2}K^{1+\kappa_{\rm D}}\right)\>\text{(SCM)}$ (23)
For OFDM, and due to the same theorem, the overhead equals
$\left(L-1+N_{\rm p,f}\right)N_{\rm p,t}=\Omega\left((\log K)^{2}K^{1+\kappa_{\rm D}}\right)\>\text{(OFDM)}$ (24)
OFDM overhead is thus smaller than its SCM counterpart but has the same asymptotic growth. AFDM overhead satisfies
[…]
As for OTFS, its overhead satisfies
[…]
Therefore, under delay-Doppler sparsity, AFDM overhead is asymptotically dominated by that of OFDM, SCM and OTFS.
The next section is dedicated to relaxing the on-grid approximation to generalize the above results to more realistic propagation and channel prediction.
IV Estimation and extrapolation of doubly sparse LTV channels with off-grid Doppler shifts
While the on-grid approximation of Definition 1 is useful for the conception of LTV channel estimation and sensing schemes and for the analysis of their performance, as we argued in Section III, it lacks support for the finer Doppler resolution needed for channel prediction or the mitigation of channel aging. Indeed, most channel prediction paradigms [28, 29, 30, 31] involve, explicitly or implicitly, the estimation of Doppler frequency shifts to within an error margin smaller than the frequency resolution characteristic of the duration of the channel observation interval. For that sake, we now present a second approximation for LTV channels that allows, in contrast to the first approximation in Definition 1, for fractional-valued Doppler frequency shifts.
@@ IV-A Off-grid doubly sparse linear time-varying channels
We now allow for fractional-valued Doppler frequency shifts. In this case, only $\iota_{p}$ is zero, while $\kappa_{p}$ may take non-zero values sampled from a uniform distribution, i.e., $\kappa_{p}\sim\mathcal{U}([-\frac{1}{2},\frac{1}{2}])$. Define $N_{\mathrm{D},l,q}$ as the number of “sub-paths”, or “rays” using the terminology of clustered delay line (CDL) models [32], which share the same integer part of their delays and Doppler shifts while only differing in the fractional part of their Doppler shifts:
@@
Depending on the scenario, it is in principle possible to model $N_{\mathrm{D},l,q}$ either as a fixed value or as a random variable, for example, a uniform random variable drawn over the range $\llbracket 1\mkern 1.5mu..\mkern 1.5muN_{\rm D}\rrbracket$. For simplicity, we opt for the first option, that is, $N_{\mathrm{D},l,q}=N_{\rm D},\forall l,q$ for some value $N_{\rm D}$ and we impose that the following channel power normalization should be satisfied.
@@
Definition 4 (off-grid Delay-Doppler sparsity).
The complex gain $h_{l,n}$ of the LTV channel $l$-th path varies with time as
@@
for some value $N_{\rm D}$, with $I_{l,q}$ retaining the same description as in Definition 1 and with $I_{l,q},\left\{\alpha_{l,q,i},\kappa_{i}\right\}_{i=1\cdots N_{\rm D}}$ being mutually independent. The complex gains satisfy $\alpha_{l,q,i}\sim\mathcal{CN}\left(0,\sigma_{\alpha}^{2}\right)$ with $\sigma_{\alpha}^{2}$ chosen so that (28) is respected.
@@
Thanks to this relaxation, the channel model becomes compatible with CDL models. Indeed, a CDL channel is composed of rays, the delay and Doppler shifts of which have a random component around a mean value that is a function of the specific cluster to which these rays belong [32]. In our model, such rays are regrouped depending on the integer part of their normalized delay and Doppler frequency shifts, i.e., as a function of the nearest delay-Doppler grid point. The only difference is that, in our model, we allow for more randomness by introducing the random variables $I_{l,q}$ which determine which points of the delay-Doppler grid will be activated (while this activation is pre-determined in CDL models by the tables defining the mean delay and Doppler shifts characterizing each cluster of the model).
Remark 3.
Maintaining the delay taps as integer values is not detrimental to the core message of this work, since the same approach we develop to deal with off-grid Doppler shifts can be extended to off-grid delays. Moreover, the effect of off-grid delays can still be made to fall under the current model by increasing the value of the model delay domain sparsity parameter $p_{\rm d}$ sufficiently to account for leakage due to the fractional part of the delay shifts.
@@
Estimating all the unknown parameters of the off-grid channel model in (29), as in super-resolution approaches such as SBL and atomic-norm minimization [12, 6, 13] or even just trying to identify them to the closest point of a finer delay-Doppler grid, as in refined-grid techniques [9, 10], remains a challenging task. Indeed, the $N_{\rm D}$ Doppler frequency shifts $\{q+\kappa_{i}\}_{i=1\cdots N_{\rm D}}$ associated with a grid point $(l,q)$ differ only by their fractional parts, which renders the problem of estimating the corresponding complex coefficients $\{\alpha_{l,q,i}\}_{i=1\cdots N_{\rm D}}$ ill-conditioned. Moreover, the number of unknowns in the model, i.e., $N_{\rm p}$, could be prohibitively large. While grid refinement methods alleviate the dimensionality problem by constraining the sub-paths to a refined grid of a controlled size, they still suffer from ill-conditioning. For both these reasons, we propose a new model based on multiple “elementary” basis expansion models (BEM), each set with a bandwidth equal to the frequency resolution and shifted in frequency to be centered at one of the active grid points. Thanks to its optimality in terms of time-frequency localization [33], we opt for discrete prolate spheroidal sequences (DPSS) BEMs.
@@ IV-B Background: DPSS Basis Expansion Model
We begin with an overview of DPSS BEM modeling of a baseband signal $h_{n}$ that occupies a digital-frequency bandwidth (W,W)(-W,W) with $W\in\left(0,\frac{1}{2}\right)$. DPSS basis vectors $u_{b,n}^{(N,W)}$ ($b=1,\ldots,N$) are the eigenvectors of the prolate matrix [34]:
$\sum_{k=0}^{N-1}C_{k,n}^{(N,W)}u_{b,k}^{(N,W)}=\lambda_{b}^{(N,W)}u_{b,n}^{(N,W)},n=0\cdots N-1$ (30)
where $C_{k,n}^{(N,W)}$ is the (kn)(k-n)-th entry of the prolate matrix:
@@
The eigenvectors are normalized so that $\sum_{n=1}^{N}\left(u_{b,n}^{(N,W)}\right)^{2}=1$. The eigenvalues are indexed in a decreasing order, $1\geq\lambda_{0}^{(N,W)}\geq\ldots\geq\lambda_{N-1}^{(N,W)}\geq 0$, and display an energy-concentration property as $\lambda_{b}^{(N,W)}$ are clustered near $1$ for $b\leq 2WN$, and rapidly drop to zero for $b>2WN$ [35]. In what follows, we will employ multiple “elementary” DPSS BEMs, each of which is used to represent the channel signal component related to the $N_{\rm D}$ fractional Doppler shifts around one of the grid points $(l,q)$, i.e., $\sum_{i=1}^{N_{\rm D}}\alpha_{l,q,i}e^{\imath 2\pi\frac{n\kappa_{i}}{N}}$, and not the channel signal associated with the whole Doppler spread, i.e., $h_{l,n}$. Each of these BEMs is defined using (30) and (31) with $W=\frac{1}{2N}$. In the interest of simplicity, all parameters with the superscript $(N,W)$ in their notation, such as $u_{b,n}^{(N,W)}$, will be replaced by their simplified forms e.g., $u_{b,n}$.
IV-C Off-grid modeling using multiple shifted elementary BEMs
The complex gain $h_{l,n}$ of the $l$-th LTV channel path, as defined in (29), is the sum of $2Q+1$ terms, each representing a signal occupying a digital-frequency band centered at $\frac{q}{N}$ and a bandwidth equal to $\frac{1}{N}$:
$h_{l,q,n}\triangleq\sum_{i=1}^{N_{\rm D}}\alpha_{l,q,i}e^{\imath 2\pi\frac{n(q+\kappa_{i})}{N}}.$ (32)
The baseband version of $h_{l,q,n}$, defined as $e^{-\imath 2\pi\frac{nq}{N}}h_{l,q,n}$, can be modeled (see Figure 7) using a DPSS BEM of an order equal to $Q_{\rm BEM}$ (chosen large enough depending on the required modeling precision) and a bandwidth (W,W)(-W,W) with $W=\frac{1}{2N}$ by defining
[…]
(a) stems: off-grid Doppler shifts, $\tilde{u}_{b,k}$: the DFT of $u_{b,n}$
(b) $N=2048$, $W=\frac{1}{2N}$, $Q=7$, $S_{\rm D}=3$
Figure 7: (a) The use of $S_{\rm D}$ frequency shifted copies of an elementary DPSS BEM $\left\{u_{b,n}\right\}_{b=1\cdots Q_{\rm BEM}}$ to capture the leakage due to $S_{\rm D}$ clusters of off-grid Doppler shifts $\left\{q_{s}+\kappa_{i}\right\}_{\begin{subarray}{c}s=1\cdots S_{\rm D},\\ i=1\cdots N_{\rm D}\end{subarray}}$ of an LTV channel tap $h_{l,n}$ (b) Precision of this representation
where the complex exponential $e^{\imath 2\pi\frac{nq}{N}}$ is used to shift, in frequency, the BEM term back to the digital frequency $\frac{q}{N}$. In vector form
@@ IV-C Off-grid modeling using multiple shifted elementary BEMs
with $\mathbf{U}_{Q_{\rm BEM}}\triangleq\left[\mathbf{u}_{1}\ldots\mathbf{u}_{Q_{\rm BEM}}\right]$ and $\mathbf{P}^{\rm BEM}=\mathbf{U}_{Q_{\rm BEM}}\mathbf{U}_{Q_{\rm BEM}}^{\rm H}$ being the orthogonal projection matrix associated with the DPSS BEM and $\mathbf{E}_{f}\triangleq\mathrm{diag}\left(e^{\imath 2\pi f0}\ldots e^{\imath 2\pi f(N-1)}\right)$. Inserting the shifted elementary BEMs, defined in (33) for each Doppler grid point, into (29) gives
@@
The precision of the above in relation to the order of its elementary BEMs is provided by the following theorem.
@@
For any $\epsilon>0$, if we set $Q_{\rm BEM}\geq C\log\frac{1}{\epsilon}$ for some constant $C$ then the error of the model in (36) satisfies $\sum_{l=0}^{L-1}\mathbb{E}[|h_{l,n}-h_{l,n}^{\rm BEM}|^{2}]<\epsilon$ for a sufficiently large $N$.
@@
Theorem 3 states that the number of DPSS basis functions needed to represent the channel component associated with a single delay-Doppler grid point grows only logarithmically with the inverse of the target precision. The following figure (as well as the numerical results of Section V) shows that as few as 4 DPSS basis functions ($Q_{\rm BEM}=4$) are sufficient to achieve good-enough precision. Note that this value is unrelated to the Doppler spread value $2Q+1$: a larger Doppler spread does not imply the need for a larger value of $Q_{\rm BEM}$.
IV-D DS-LTV off-grid LMMSE channel estimation
After relaxing the integer-Doppler assumption, thanks to the “multiple shifted elementary BEMs” model, we now undertake another step towards the goal of practical channel estimation and prediction for sparse LTV channels. Indeed, while the compressed-sensing paradigm provided us with theoretical tools that helped in the analytical comparison of the performance of different waveforms, it is preferable to perform channel estimation using LMMSE due to its computational-complexity advantage. Let $\boldsymbol{\beta}$ be the vectorized form of the BEM coefficients $\{\beta_{l,q,b}\}_{\begin{subarray}{c}b=1\cdots Q_{\rm BEM}\\ l,q,I_{l,q}\neq 0\end{subarray}}$ associated with the active delay-Doppler grid points. Note that the knowledge of the delay-Doppler sparsity support, i.e., of $\left\{I_{l,q}\right\}_{l=0\cdots L-1,q=-Q\cdots Q}$, at the receiver side is now assumed. Define $\mathbf{B}$ as the following block-diagonal matrix
[…]
where $\widetilde{\mathbf{B}}\triangleq[\widetilde{\mathbf{B}}_{-Q}\ldots\widetilde{\mathbf{B}}_{Q}]$ is a $N\times(2Q+1)Q_{\rm BEM}$ matrix, $\widetilde{\mathbf{B}}_{q}]_{n,b}=\frac{1}{\sqrt{N}}\tilde{U}_{b,(n-q)}$ and $\tilde{U}_{b,n}=\frac{1}{\sqrt{N}}\sum_{k=0}^{N-1}u_{b,k}e^{-\imath 2\pi\frac{nk}{N}}$ is the DFT of the DPSS basis vector $u_{b,n}$. Let $\mathbf{A}_{\beta}$ be the matrix that places the blocks of $\boldsymbol{\beta}$, each of size $Q_{\rm BEM}$, within the positions corresponding to $I_{l,q}\neq 0$ in a larger vector of length $LQ_{\rm BEM}(2Q+1)$ resulting in a $LQ_{\rm BEM}(2Q+1)$-long vector $\mathbf{A}_{\beta}\boldsymbol{\beta}$ that is block sparse. Now, define